Using matrix calculations, solve the following problem. Show Your Work.
A total of $50,000 is invested in three funds paying 6%, 8%, and 10% simple interest. The yearly interest is $3,700. Twice as much money is invested at 6% as invested at 10%. How much is invested in each of the funds.

set up the 3 linear equations in 3 variables
put the coefficients into a matrix
solve the matrix using any method you like
provide the solutions for x, y, and z.
I have two equations already which are:
.06x+.08y+.10z=3,700
x+y+z=50,000
How and where do I get the last equation? I am not really sure what to do with this problem from here. Can someone please help me. I have been at this all day. Thank you

Answer by stanbon(48569)

Using matrix calculations, solve the following problem. Show Your Work.
A total of $50,000 is invested in three funds paying 6%, 8%, and 10% simple interest. The yearly interest is $3,700. Twice as much money is invested at 6% as invested at 10%. How much is invested in each of the funds.
set up the 3 linear equations in 3 variables
put the coefficients into a matrix
solve the matrix using any method you like
provide the solutions for x, y, and z
----------------------------
Quantity Equation: x + y + z = 50,000
Interest EQUATION:0.06x + 0.08y + 0.10z = 3700
Relation Equation: x + 0 - 2z = 0
--------------------------
Using the Matrix function of a TI calculator I got:
x = 30000
y = 5000
z = 15000
===================
Cheers,
Stan H.